PULSED MOLECULAR BEAM FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF α-PINENE AND β-PINENE

\underline{DALE J. BRUGH} and RICHARD D. SUENRAM, National Institute of Standards and Technology, Optical Technology Division, Gaithersburg, MD 20899-8441.

Chemical analysis of fire debris is often conducted to determine whether accelerants were employed. The GC and GC/MS methods used in these analyses do not always offer the chemical specificity that the forensic chemist would like, however. In an effort to demonstrate the potential application of the Fourier-transform microwave (FTMW) spectrometer as a chemically-specific forensic tool we have collected the pure rotational spectra of the major constituents of one possible accelerant, turpentine. The major constituents are α-pinene and β-pinene, and their spectra have been recorded from 8 GHz to 20 GHz. The spectrum of β-pinene has been assigned, and the rotational constants A = 1901.8889 ± 0.00048 MHz, B = 1293.66095 ± 0.00026 MHz, and C = 1150.83126 ± 0.00010 MHz, as well as several distortion constants, have been extracted from the data. Stark data have also been collected, from which the dipole moment will be determined. The spectrum of α-pinene is currently under analysis. Because the rotational spectra for α-pinene and β-pinene are distinct, the relative amounts of the two constituents in a sample could be determined, something that would be useful in deciding if the detected pinenes are of natural orgin or not.